There is a set of twins, one an


astronaut, the other works for mission control of NASA. The astronaut leaves on


a deep space trip traveling at 95% the speed of light. Upon returning the


astronauts clock has measured ten years, so the astronaut has aged 10


years. What is the time that the


earth-bound brother measures while his twin was on his exploration in space?


How old is he?

1
Expert's answer
2022-02-02T17:34:36-0500

"t=\frac {t_0}{\sqrt{1-\frac{v^2}{c^2}}}=32" years.